# Schrödinger Bridge Matching for Tree-Structured Costs and Entropic Wasserstein Barycentres

### v1

@@ 5 Discussion
Beyond the IPF and IMF approaches already mentioned, other approaches for the SB problem include adversarial solvers (Kim et al., 2024; Gushchin et al., 2023), parameterisation via potentials for Gaussian mixtures (Korotin et al., 2024; Gushchin et al., 2024), and variational approaches (Deng et al., 2024). For the Wasserstein barycentre problem, standard approaches ensure tractability by representing the barycentre with a finite set of points (either updating only weightings (Benamou et al., 2015; Solomon et al., 2015; Cuturi and Peyré, 2016; Staib et al., 2017; Dvurechenskii et al., 2018), or also the positions of the points in the support (Rabin et al., 2012; Cuturi and Doucet, 2014; Claici et al., 2018; Luise et al., 2019)). Such approaches are effective in lower dimensions but scale poorly, cannot be used to generate new samples, and do not capture the true continuous nature of the barycentre in continuous settings. To address these limitations, recent work has focused on learning continuous approximations of the barycentre. Li et al. (2020) optimise for the regularised dual potentials for general costs, while others parameterise Wasserstein-2 potentials with convex neural architectures using adversarial losses (Fan et al., 2021) or additional cycle-consistency regularisation (Korotin et al., 2021). Korotin et al. (2022) leverage the fixed-point property in Álvarez Esteban et al. (2016). We also highlight a recent line of works Kolesov et al. (2024b), Kolesov et al. (2024a), Gazdieva et al. (2025) which consider bi-level adversarial approaches to approximate the barycentre for general cost functions. Recent work Visentin and Cheridito (2025) solves for the barycentre by parametrising transport maps using normalising flows, and performing gradient descent on the total transport cost.
Our approach utilises non-adversarial bridge-matching loss objectives, which provide stable training. This is at a cost of requiring multiple sequential IMF iterations. This iterative nature resembles fixed-point barycentre methods, however our approach differs in that each iteration requires only performing bridge-matching rather than solving an entire OT problem along each edge. Finally, we emphasise that our approach tackles the tree-structured problem considered in Haasler et al. (2021); Noble et al. (2023), and thus is applicable beyond only barycentre problems. We provide an an example applying TreeDSBM to a non-star-shaped tree in Appendix E.
@@
Our method shares the same limitations as other flow-based SB solvers. It is restricted to quadratic cost functions for OT, and introduces an entropic bias. Inference is expensive in comparison to methods that use a single function evaluation, and our method is not simulation-free, requiring simulations of the current learned processes at each iteration. Future advances in the flow and SB literatures will aid in addressing these limitations.
@@
We have extended the IMF procedure to the tree-structured SB problem, providing a scalable flow-based approach that comes with a well-understood theoretical analysis, and in particular can be used for entropic Wasserstein barycentre computation. Our TreeDSBM algorithm displays improved performance over its IPF counterpart TreeDSB and is competitive with current state-of-the-art continuous Wasserstein-2 barycentre solvers, demonstrating that flow-based approaches for barycentre estimation can offer a compelling alternative to established methods. Future directions can investigate improved architectures and implementation techniques inspired by progress in the flow-matching literature, as well as extensions to other data modalities.
@@ Acknowledgements
SH and PP are supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning [grant number EP/S023151/1]. GD was supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
@@
The Appendix is structured in the following manner. In Appendix A, we provide further background information regarding the problems discussed in the paper. In Appendix B, we provide the proofs of the results stated in the main body. In Appendix C, we describe in more detail how to implement the TreeDSBM algorithm, and also discuss extensions of the methodology for improving convergence speed. In Appendix D, we give a full overview of the experimental setups in the main body, and include further discussions of the findings. In Appendix E we provide additional experiments not included in the main body. Finally, Appendix F includes the licenses for assets used in this work.
@@ Appendix A Background
A.1 Optimal Transport
We here provide a brief overview of optimal transport; for more details we refer to Santambrogio [2015], Peyré and Cuturi [2019].
@@ A.1 Optimal Transport
For two measures $\mu_{0}$ and $\mu_{1}$, the original OT formulation given by Monge [1781] aims to find a map $T^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ pushing $\mu_{0}$ onto $\mu_{1}$, that is $\mu_{1}=T^{*}\#\mu_{0}$, while minimising the total cost of transportation according to a given cost function $c:{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$,
@@
While intuitive, the Monge formulation of OT is restrictive in that it does not permit splitting of mass in the transportation. Instead, it is common to consider the more general Kantorovich formulation [Kantorovich, 1942], which searches over joint distributions $\pi\in\mathcal{P}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$ with the correct marginals,
@@
The resulting problem has many desirable properties; it enables differentiability with respect to the inputs, relaxes constraints on the corresponding potentials [Genevay et al., 2016], and for discrete measures enables efficient computation using Sinkhorn’s algorithm [Cuturi, 2013].
A.2 Multi-Marginal Optimal Transport
One can generalise the above optimal transport problem to the case of multiple marginal distributions. For a cost function $c:({\mathbb{R}}^{d})^{\ell+1}\rightarrow{\mathbb{R}}$ and an ‘observed’ set $\mathcal{S}\subset\{0,...,\ell\}$, multi-marginal optimal transport (mmOT) searches over joint distributions $\pi\in\mathcal{P}(({\mathbb{R}}^{d})^{\ell+1})$ matching the prescribed marginals on $\mathcal{S}$ to minimise the objective,
@@ A.2 Multi-Marginal Optimal Transport
A.3 Wasserstein Barycentres
Given $\ell$ measures $(\mu_{1},...,\mu_{\ell})$ and weights $(\lambda_{1},...,\lambda_{\ell})$ summing to 1, the Wasserstein-2 barycentre [Agueh and Carlier, 2011] is defined as
@@ A.3 Wasserstein Barycentres
where $W_{2}^{2}(\mu_{i},\nu)$ denotes the minimum attained by the OT solution in (18) for quadratic cost function $c(x,y)=\tfrac{1}{2}\lVert x-y\rVert_{2}^{2}$. Observe that this can be cast as an mmOT problem by using a star-shaped cost function $c(x_{0:\ell})=\sum_{i=1}^{\ell}\lambda_{i}\lVert x_{0}-x_{i}\rVert_{2}^{2},$ where the marginals are prescribed on the leaf nodes $\mathcal{S}=\{1,...,\ell\}$, and the centre vertex 00 is an unobserved measure (which at the solution is the barycentre). The Wasserstein barycentre is widely studied due to its importance in applications, including in Bayesian learning [Srivastava et al., 2018], clustering [Ye et al., 2017], and representation learning [Singh et al., 2020] to name a few.
@@
There are many different ways to define an entropy-regularised formulation of the Wasserstein barycentre problem. Some formulations add inner regularisation, which replaces the Wasserstein distances with an entropy-regularised version, while others also consider outer regularisation in which an entropic penalty on the barycentre is added. The $(\varepsilon,\tau)$-doubly-regularised barycentre of Chizat [2023] unifies many of these problems, and aims to minimise the objective
@@
where $W_{2,\varepsilon}^{2}(\mu_{i},\nu)=\min_{\pi:\pi_{0}=\mu_{i},\pi_{1}=\nu}\big{\{}\iint\tfrac{1}{2}\lVert x-y\rVert^{2}{\mathrm{d}}\pi(x,y)+\varepsilon{\mathrm{KL}}(\pi\|\mu\otimes\nu)\big{\}}$. We refer to Chizat [2023] for more details regarding the different types of entropy-regularised Wasserstein barycentres.
@@
The case of tree-structured costs for multi-marginal optimal transport is often specifically studied in the literature [Haasler et al., 2021, Noble et al., 2023], as it recovers Wasserstein barycentres as a special case as described above, and also arises in the Wasserstein propagation problem [Solomon et al., 2014, 2015]. In the discrete setting, Haasler et al. [2021] show that the tree structure can be leveraged to design an efficient Sinkhorn-based algorithm.
@@
Recall from Section 2.3 that we identify a tree $\mathcal{T}$ (with vertices, edge set, and length function $(\mathcal{V},\mathcal{E},\ell)$) with a metric space, by associating each edge $e$ with the interval $[0,\ell(e)]$, which are connected according to the tree structure. The same construction is used in Noble et al. [2023] for defining the dynamic TreeSB problem that we study in this work. For a rigorous description of such constructions, see for example Hambly and Lyons [2008], Bolin et al. [2024].
@@ Appendix B Proofs
In this section, we give proofs of the results stated in the main paper. We first provide the proof of the TreeSB characterisation in Theorem 3.1. We will follow the presentation and proof techniques of Léonard [2013], making the appropriate changes to extend to the tree-structured case. We then prove the convergence of the TreeIMF procedure stated in Theorem 3.1, following the presentation and proof techniques of Shi et al. [2023].
@@ B.1 Existence and Uniqueness of TreeSB solution
We first provide results pertaining to the existence and uniqueness of the TreeSB solution. Note that while we follow the presentation of Léonard [2013], one could instead rely on the existence results in Noble et al. [2023].
@@
We note here that the reference measures that we consider in the various SB problems are unbounded, as we consider the reference Brownian motions to be in stationarity. We will make use of properties of the KL divergence defined relative to these measures; see the Appendix of Léonard [2013] for a justification of why these properties still hold when the reference measures are unbounded.
@@
Proposition B.1 (Compare to Proposition 2.3 in Léonard [2013]).
[…]
@@
We see that ${\mathrm{KL}}(\mathbb{P}\|{\mathbb{Q}})\geq{\mathrm{KL}}(\mathbb{P}_{\mathcal{V}}\|{\mathbb{Q}}_{\mathcal{V}})$, with equality if and only if $\mathbb{P}(\cdot|X_{\mathcal{V}})={\mathbb{Q}}(\cdot|X_{\mathcal{V}})$ for $\mathbb{P}_{\mathcal{V}}$-a.e. $X_{\mathcal{V}}$ (assuming $KL(\mathbb{P}\|{\mathbb{Q}})<\infty$). Therefore, $\mathbb{P}$ solves the dynamic problem if and only if it decomposes as a mixture over bridges ${\mathbb{Q}}_{\cdot|X_{\mathcal{V}}}$ according to the coupling $\mathbb{P}_{\mathcal{V}}$ solving the static problem, i.e. $\mathbb{P}^{SB}=\Pi_{\mathcal{V}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{V}}$ (if this were not true, then $\mathbb{P}^{SB}_{\mathcal{V}}{\mathbb{Q}}_{\cdot|\mathcal{V}}$ would be a valid solution with lower KL divergence relative to ${\mathbb{Q}}$, contradicting optimality of $\mathbb{P}^{SB}$). Note that such bridges just consist of Brownian bridges along the individual edges.
@@
Now we have ${\mathrm{KL}}(\mathbb{P}\|{\mathbb{Q}})\geq{\mathrm{KL}}(\mathbb{P}_{\mathcal{S}}\|{\mathbb{Q}}_{\mathcal{S}})$, with equality if and only if $\mathbb{P}(\cdot|X_{\mathcal{S}})={\mathbb{Q}}(\cdot|X_{\mathcal{S}})$ for $\mathbb{P}_{\mathcal{S}}$-a.e. $X_{\mathcal{S}}$ (assuming ${\mathrm{KL}}(\mathbb{P}\|{\mathbb{Q}})<\infty$). So in particular, we have that $\mathbb{P}^{SB}$ must be a mixture of ${\mathbb{Q}}$-bridges according to its own coupling over $\mathcal{S}$, i.e. $\mathbb{P}^{SB}=\mathbb{P}_{\mathcal{S}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{S}}$ ∎
In the following, we will utilise this equivalence between the ($\mathrm{TreeSB}_{\mathrm{dyn}}$) and ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$) problems. We now present an auxiliary result giving a criterion for the tree-structured SB problems to have a solution, in the vein of Lemma 2.4 in Léonard [2013]. This is a technical result required to deal with the fact that we are considering an unbounded reference measure ${\mathbb{Q}}$.
Lemma B.2 (Compare to Lemma 2.4 in Léonard [2013]).
@@
[…]
@@
In light of the equivalence in Proposition B.1, we can consider just the static problem ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$). Since the marginals $\mu_{i}$ are tight on ${\mathbb{R}}^{d}$, it easily follows that the closed constraint set $\Gamma(\{\mu_{i}\}_{i\in\mathcal{S}}):=\{\Pi\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{S}|}:\Pi_{i}=\mu_{i}~{}~{}\forall i\in\mathcal{S}\}$ is uniformly tight and thus compact in $\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{S}|}$. From the characterisation of KL divergence with respect to the unbounded measure ${\mathbb{Q}}_{\mathcal{S}}$ (see Léonard [2013], Appendix A), we have ${\mathrm{KL}}(\Pi\|{\mathbb{Q}}_{\mathcal{S}})={\mathrm{KL}}(\Pi\|{\mathbb{Q}}_{\mathcal{S}}^{B})-\int_{({\mathbb{R}}^{d})^{|\mathcal{S}|}}\sum_{i\in\mathcal{S}}B(x_{i}){\mathrm{d}}\Pi(x_{\mathcal{S}})-z_{B}$, where ${\mathbb{Q}}_{\mathcal{S}}^{B}$ is the normalised measure ${\mathbb{Q}}_{\mathcal{S}}^{B}=\tfrac{1}{z_{B}}\exp(-\bigoplus_{\mathcal{S}}B){\mathbb{Q}}_{\mathcal{S}}$. For $\Pi\in\Gamma(\{\mu_{i}\}_{i\in\mathcal{S}})$, we have
@@
Proposition B.3 (Compare to Proposition 2.5 in Léonard [2013]).

### v2

@@ 5 Discussion
Beyond the IPF and IMF approaches, other approaches for the SB problem include adversarial solvers [43, 37], parameterisation via potentials for Gaussian mixtures [48, 36], and variational approaches [25]. For Wasserstein barycentre computation, standard approaches ensure tractability by representing the solution with a finite set of points (either updating only weightings [6, 79, 20, 85, 27], or also the positions of the points in the support [71, 18, 15, 59]). Such approaches are effective in lower dimensions but scale poorly, cannot be used to generate new samples, and do not capture the true continuous nature of the barycentre. To address these limitations, recent work has focused on learning continuous approximations. [54] optimise for the regularised dual potentials for general costs, while others parameterise Wasserstein-2 potentials with convex neural architectures using adversarial losses [28] or additional cycle-consistency regularisation [49]. [47] leverage the fixed-point property in [4]. We also highlight a recent line of works of [46, 45] which consider bi-level adversarial approaches for general cost functions. Our approach utilises non-adversarial bridge-matching loss objectives, which provide stable training but requires multiple sequential IMF iterations. Finally, we emphasise that our approach tackles the tree-structured SB problem [38, 64], and thus is applicable beyond only barycentre problems (we consider a toy example in Appendix E.3).
@@
Our method shares the same limitations as other flow-based SB solvers. It is restricted to quadratic cost functions for OT, and introduces an entropic bias. Inference is expensive in comparison to methods that use a single function evaluation, and our method is not sampling-free, requiring simulations of the current learned processes at each iteration. Future advances in the flow and SB literatures can aid in addressing these limitations. We also provide an additional experiment in Appendix E that highlights a possible limitation of our method in scenarios where there is a simple shared structure between the known marginals and the barycentre.
@@
We have extended the IMF procedure to the tree-structure SB problem, providing a scalable flow-based approach that in particular can be used for entropic Wasserstein barycentre computation. Our TreeDSBM algorithm displays improved performance over its IPF counterpart TreeDSB, and demonstrates that flow-based approaches for barycentre estimation can offer a compelling alternative to established continuous Wasserstein-2 barycentre algorithms. Future directions can investigate improved architectures and implementation techniques inspired by progress in the flow-matching literature, as well as extensions to other data modalities.
@@ Acknowledgements
SH is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning [grant number EP/S023151/1]. PP is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning [EP/S023151/1], a Google PhD Fellowship, and an NSERC Postgraduate Scholarship (PGS D). GD was supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1]. The authors would like to thank James Thornton for helpful discussions.
@@
The Appendix is structured in the following way. In Appendix A, we define the notation used and provide further background information regarding the problems discussed in the paper. In Appendix B, we provide the proofs of the results stated in the main body. In Appendix C, we describe in more detail how to implement the TreeDSBM algorithm, and also discuss extensions of the methodology for improving convergence speed. In Appendix D, we give a full overview of the experimental setups in the main body, and include further discussions of the findings. In Appendix E we provide additional experiments that could not be included in the main body due to space constraints. Finally, Appendix F includes the licenses for assets used in this work.
@@ Appendix A Background
A.1 Notation
[…]
A.2 Optimal transport
We here provide a brief overview of optimal transport; for more details we refer to [74], [69].
@@ A.2 Optimal transport
For two measures $\mu_{0}$ and $\mu_{1}$, the original OT formulation given by [63] aims to find a map $T^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ pushing $\mu_{0}$ onto $\mu_{1}$, that is $\mu_{1}=T^{*}\#\mu_{0}$, while minimising the total cost of transportation according to a given cost function $c:{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$,
@@
While intuitive, the Monge formulation of OT is restrictive in that it does not permit splitting of mass in the transportation. Instead, it is common to consider the more general Kantorovich formulation [40], which searches over joint distributions $\pi\in\mathcal{P}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$ with the correct marginals,
@@
The resulting problem has many desirable properties; it enables differentiability with respect to the inputs, relaxes constraints on the corresponding potentials [34], and for discrete measures enables efficient computation using Sinkhorn’s algorithm [17].
A.3 Multi-marginal optimal transport
One can generalise the above optimal transport problem to the case of multiple marginal distributions. For a cost function $c:({\mathbb{R}}^{d})^{\ell+1}\rightarrow{\mathbb{R}}$ and an ‘observed’ set $\mathcal{S}\subset\{0,$$\ell\}$, multi-marginal optimal transport (mmOT) searches over joint distributions $\pi\in\mathcal{P}(({\mathbb{R}}^{d})^{\ell+1})$ matching the prescribed marginals on $\mathcal{S}$ to minimise the objective,
@@ A.3 Multi-marginal optimal transport
Multi-marginal Schrödinger bridges
We emphasise here the distinction between the multi-marginal OT and corresponding SB problem that we consider in this work, compared to the multi-marginal SB problem that has been considered recently in works such as [13, 11, 76] to name a few. These lines of works aim to find processes that evolve in time while fitting to known marginals at several different timepoints, and are motivated by inferring population dynamics from snapshot data. We rather consider costs defined according to a tree structure, in which the marginals at some of the vertices may not be known (as considered in [38, 64]), with a particular focus on applications to Wasserstein barycentre problems.
A.4 Wasserstein barycentres
Given $\ell$ measures $(\mu_{1},...,\mu_{\ell})$ and weights $(\lambda_{1},...,\lambda_{\ell})$ summing to 1, the Wasserstein-2 barycentre [1] is defined as
@@ A.4 Wasserstein barycentres
where $W_{2}^{2}(\mu_{i},\nu)$ denotes the minimum attained by the OT solution in (17) for quadratic cost function $c(x,y)=\tfrac{1}{2}\lVert x-y\rVert_{2}^{2}$. Observe that this can be cast as an mmOT problem by using a star-shaped cost function $c(x_{0:\ell})=\sum_{i=1}^{\ell}\lambda_{i}\lVert x_{0}-x_{i}\rVert_{2}^{2},$ where the marginals are prescribed on the leaf nodes $\mathcal{S}=\{1,$$\ell\}$ and the centre vertex 0 is an unobserved measure (which at the solution is the barycentre). The Wasserstein barycentre is widely studied due to its importance in applications, including in Bayesian learning [84], clustering [90], and representation learning [78] to name a few.
@@
There are many different ways to define an entropy-regularised formulation of the Wasserstein barycentre problem. Some formulations add inner regularisation, which replaces the Wasserstein distances with an entropy-regularised version, while others also consider outer regularisation in which an entropic penalty on the barycentre is added. The $(\varepsilon,\tau)$-doubly-regularised barycentre of [14] unifies many of these problems, and aims to minimise the objective
@@
where $W_{2,\varepsilon}^{2}(\mu_{i},\nu)=\min_{\pi:\pi_{0}=\mu_{i},\pi_{1}=\nu}\big\{\iint\tfrac{1}{2}\lVert x-y\rVert^{2}{\mathrm{d}}\pi(x,y)+\varepsilon{\mathrm{KL}}(\pi\|\mu\otimes\nu)\big\}$. We refer to [14] for more details regarding the different types of entropy-regularised Wasserstein barycentres.
@@
The case of tree-structured costs for multi-marginal optimal transport is often specifically studied in the literature [38, 64], as it recovers Wasserstein barycentres as a special case as described above, and also arises in the Wasserstein propagation problem [80, 79]. In the discrete setting, [38] show that the tree structure can be leveraged to design an efficient Sinkhorn-based algorithm.
@@
Recall from Section 2.3 that we identify a tree $\mathcal{T}$ (with vertices, edge set, and length function $(\mathcal{V},\mathcal{E},\ell)$) with a metric space, by associating each edge $e$ with the interval $0,\ell(e)]$, which are connected according to the tree structure. The same construction is used in [64] for defining the dynamic TreeSB problem that we study in this work. For a rigorous description of such constructions, see for example [39], [8].
@@ Appendix B Proofs
In this section, we give proofs of the results stated in the main paper. We first provide the proof of the TreeSB characterisation in Theorem 3.1. We will follow the presentation and proof techniques of [60], making the appropriate changes to extend to the tree-structured case. We then prove the convergence of the TreeIMF procedure stated in Theorem 3.1, following the presentation and proof techniques of [77].
@@ B.1 Existence and uniqueness of TreeSB solution
We first provide results pertaining to the existence and uniqueness of the TreeSB solution. Note that while we follow the presentation of [60], one could instead rely on the existence results in [64].
@@
We note here that the reference measures that we consider in the various SB problems are unbounded, as we consider the reference Brownian motions to be in stationarity. We will make use of properties of the KL divergence defined relative to these measures; see the Appendix of [60] for a justification of why these properties still hold when the reference measures are unbounded.
@@
Proposition B.1 (Compare to Proposition 2.3 in [60]).
The tree-structured Schrödinger Bridge problems ($\mathrm{TreeSB}_{\mathrm{dyn}}$), ($\mathrm{TreeSB}_{\mathrm{stat}}$), and ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$) admit at most one solution $\mathbb{P}^{SB}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}})$}, $\Pi_{\mathcal{V}}^{SB}\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{V}|})$, and $\Pi_{\mathcal{S}}^{SB}\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{S}|})$ respectively.
If $\mathbb{P}^{SB}$ solves ($\mathrm{TreeSB}_{\mathrm{dyn}}$), then $\mathbb{P}^{SB}_{\mathcal{V}}$ solves ($\mathrm{TreeSB}_{\mathrm{stat}}$). Conversely, if $\Pi_{\mathcal{V}}^{SB}$ solves ($\mathrm{TreeSB}_{\mathrm{stat}}$), then ($\mathrm{TreeSB}_{\mathrm{dyn}}$) is solved by mixing Brownian bridges along each edge as $\mathbb{P}^{SB}=\Pi_{\mathcal{V}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{V}}$.
Moreover, if $\mathbb{P}^{SB}$ solves ($\mathrm{TreeSB}_{\mathrm{dyn}}$), then $\mathbb{P}^{SB}_{\mathcal{S}}$ solves ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$). Conversely, if $\Pi_{\mathcal{S}}^{SB}$ solves ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$), then ($\mathrm{TreeSB}_{\mathrm{dyn}}$) is solved by the corresponding mixture of $\mathbb{Q}$}-bridges conditioned on values at $\mathcal{S}$, $\mathbb{P}^{SB}=\Pi_{\mathcal{S}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{S}}$.
@@
We see that ${\mathrm{KL}}(\mathbb{P}\|{\mathbb{Q}})\geq{\mathrm{KL}}(\mathbb{P}_{\mathcal{V}}\|{\mathbb{Q}}_{\mathcal{V}})$, with equality if and only if $\mathbb{P}(\cdot|X_{\mathcal{V}})={\mathbb{Q}}(\cdot|X_{\mathcal{V}})$ for $\mathbb{P}_{\mathcal{V}}$-a.e. $X_{\mathcal{V}}$ (assuming $KL(\mathbb{P}\|{\mathbb{Q}})<\infty$). Therefore, $\mathbb{P}$ solves the dynamic problem if and only if it decomposes as a mixture over bridges ${\mathbb{Q}}_{\cdot|\mathcal{V}}$ according to the coupling $\mathbb{P}_{\mathcal{V}}$ solving the static problem, i.e. $\mathbb{P}^{SB}=\Pi_{\mathcal{V}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{V}}$ (if this were not true, then $\mathbb{P}^{SB}_{\mathcal{V}}{\mathbb{Q}}_{\cdot|\mathcal{V}}$ would be a valid solution with lower KL divergence relative to $\mathbb{Q}$}, contradicting optimality of $\mathbb{P}^{SB}$). Note that such bridges just consist of Brownian bridges along the individual edges.
@@
Now we have ${\mathrm{KL}}(\mathbb{P}\|{\mathbb{Q}})\geq{\mathrm{KL}}(\mathbb{P}_{\mathcal{S}}\|{\mathbb{Q}}_{\mathcal{S}})$, with equality if and only if $\mathbb{P}(\cdot|X_{\mathcal{S}})={\mathbb{Q}}(\cdot|X_{\mathcal{S}})$ for $\mathbb{P}_{\mathcal{S}}$-a.e. $X_{\mathcal{S}}$ (assuming ${\mathrm{KL}}(\mathbb{P}\|{\mathbb{Q}})<\infty$). So in particular, we have that $\mathbb{P}^{SB}$ must be a mixture of $\mathbb{Q}$}-bridges according to its own coupling over $\mathcal{S}$, i.e. $\mathbb{P}^{SB}=\mathbb{P}_{\mathcal{S}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{S}}$ . ∎
In the following, we will utilise this equivalence between the ($\mathrm{TreeSB}_{\mathrm{dyn}}$) and ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$) problems. We now present an auxiliary result giving a criterion for the tree-structured SB problems to have a solution, in the vein of Lemma 2.4 in [60]. This is a technical result required to deal with the fact that we are considering an unbounded reference measure $\mathbb{Q}$}.
Lemma B.2 (Compare to Lemma 2.4 in [60]).
@@
Note that $\inf($$\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$$)=\inf($$\mathrm{TreeSB}_{\mathrm{dyn}}$$)\in(-\infty,\infty]$ (from the previous result). The static and dynamic tree-structured SB problems ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$) and ($\mathrm{TreeSB}_{\mathrm{dyn}}$) for the $\mu_{i}$ have a (unique) solution if and only if $\inf($$\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$$)=\inf($$\mathrm{TreeSB}_{\mathrm{dyn}}$$)<\infty$ (that is, if and only if the marginals $\mu_{i}$ are such that there exists some $\Pi^{0}\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{S}|})$ satisfying $\Pi_{i}^{0}=\mu_{i}$ for each $i\in\mathcal{S}$, and ${\mathrm{KL}}(\Pi^{0}\|{\mathbb{Q}}_{\mathcal{S}})<\infty)$.
@@
In light of the equivalence in Proposition B.1, we can consider just the static problem ($\mathrm{TreeSB}_{\mathrm{stat}}^{\mathcal{S}}$). Since the marginals $\mu_{i}$ are tight on ${\mathbb{R}}^{d}$, it easily follows that the closed constraint set $\Gamma(\{\mu_{i}\}_{i\in\mathcal{S}}):=\{\Pi\in\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{S}|}:\Pi_{i}=\mu_{i}~~\forall i\in\mathcal{S}\}$ is uniformly tight and thus compact in $\mathcal{P}(({\mathbb{R}}^{d})^{|\mathcal{S}|}$. From the characterisation of KL divergence with respect to the unbounded measure ${\mathbb{Q}}_{\mathcal{S}}$ (see [60], Appendix A), we have ${\mathrm{KL}}(\Pi\|{\mathbb{Q}}_{\mathcal{S}})={\mathrm{KL}}(\Pi\|{\mathbb{Q}}_{\mathcal{S}}^{B})-\int_{({\mathbb{R}}^{d})^{|\mathcal{S}|}}\sum_{i\in\mathcal{S}}B(x_{i}){\mathrm{d}}\Pi(x_{\mathcal{S}})-z_{B}$, where ${\mathbb{Q}}_{\mathcal{S}}^{B}$ is the normalised measure ${\mathbb{Q}}_{\mathcal{S}}^{B}=\tfrac{1}{z_{B}}\exp(-\bigoplus_{\mathcal{S}}B){\mathbb{Q}}_{\mathcal{S}}$. For $\Pi\in\Gamma(\{\mu_{i}\}_{i\in\mathcal{S}})$, we have
@@
Proposition B.3 (Compare to Proposition 2.5 in [60]).